FIRST RESULTS FROM THE FAR INFRARED BEAMLINE AT THE CANADIAN LIGHT SOURCE: HIGH RESOLUTION ANALYSIS OF ACROLEIN IN THE 600 CM^{−1} REGION

\underline{A.R.W. McKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON K1A 0R6, Canada; D.W. TOKARYK, Department of Physics and Centre for Laser, Atomic, and Molecular Sciences, University of New Brunswick, Fredericton, NB E3B 5A3, Canada; LI-HONG XU, Department of Physical Sciences and Centre for Laser, Atomic, and Molecular Sciences, University of New Brunswick, Saint John, NB E2L 4L5, Canada; D.R.T. APPADOO and T. MAY, Canadian Light Source, 101 Perimeter Road, University of Saskatchewan, Saskatoon, SK S7N 0X4, Canada.

Synchrotron radiation from the new Canadian Light Source facility has been used to obtain a high resolution (0.0012 cm^{−1}) absorption spectrum of acrolein, CH_{2}CHCHO, in the 550-660 cm^{−1}region. Almost 2000 transitions have been included in a detailed analysis of the ν_{12} (∼564 cm^{−1}) and ν_{17} (∼593 cm^{−1}) fundamental bands to obtain precise values for the band origins, rotational and centrifugal distortion parameters. The analysis included the a- and b-type Coriolis interactions connecting ν_{12} and ν_{17}, as well as an a-type Coriolis interaction between ν_{17} and a “dark” perturbing state, identified as 4ν_{18}. We believe that this is the first high resolution infrared study of acrolein.